AOB is a triangle with coordinates of A and O as (4,0) and (0,0),AB=5.Find the coordinates of B.
1
Expert's answer
2012-09-04T10:04:46-0400
In common, (x(B) - 4)^2 + (y(B) - 0)^2 = AB^2.
BO^2 + AO^2 = AB^2
AO =
sqrt( (0 - 0)^2 + (4 - 0)^2 ) = 4
AB = 5
BO = sqrt( AB^2 - AO^2 ) =
3
x(B) = x(0) = 0
y(B) = sqrt( BO^2 - (x(B))^2 ) = 3
So, getting
B(0,3).